If a passenger train travels two times as fast as a freight train, and freight train takes 3 hours longer to travel 210 miles, what is the speed of each train?
Answer by TimothyLamb(4379)

x = speed of passenger train
y = speed of freight train
x = 2y
---
s = d/t
t = d/s
---
passenger train:
t = 210/x
t = 210/2y
---
freight train:
t + 3 = 210/y
t = 210/y - 3
---
210/2y = 210/y - 3
105/y = 210/y - 3
210/y - 105/y = 3
105/y = 3
y = 105/3
y = 35
x = 70
---
answer:
x = speed of passenger train = 70 mph
y = speed of freight train = 35 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php